# Biologically-Inspired Technologies: Integrating Brain-Computer Interface and Neuromorphic Computing for Human Digital Twins

### v2

@@ IV-C Case Study
By introducing the SNN-driven FL for Bio-HDT, the potential challenges aforementioned are efficiently addressed. We demonstrate the efficacy of our framework through comparison simulations in this section.
As aforementioned, the Bio-HDT framework is primarily designed to enhance three key factors: personalization, energy efficiency, and privacy preservation. In the human-centric Metaverse, personalization relies heavily on the accurate interpretation of human’s behavior intention and states, ensuring that user interactions are not only immersive but also highly responsive. For MSPs, the ability to accurately identify and interpret these signals is crucial. This capability is fundamental to effectively analyzing user requirements and creating personalized services that meet individual needs. The precision in processing these signals directly influences the quality, relevance, and timeliness of the content delivered. Consequently, it impacts both the QoS provided by the MSPs and the QoE perceived by the users. Additionally, energy efficiency is crucial for enabling sustainable and scalable operations in real-time environments, ensuring that the system can operate continuously without excessive resource consumption. Therefore, we evaluate our scheme’s performance based on the identification accuracy of brain signals and the energy consumption of neural network operations. Note that privacy is not evaluated in this case study because our SNN-driven FL framework effectively addresses privacy concerns by processing data locally. The settings of simulation are as follows:
IV-C1 Bio-HDT setting
@@ IV-C3 FL training and baselines
We utilize the FedAvg algorithm [14] to conduct the FL framework. For the learning hyper-parameters, the learning rate, batch size and local epoch are set as $\left\{0.01,64,1\right\}$. Moreover, we conduct performance comparisons between our proposed method, denoted as FL4BCI-SNN, and two baseline methods: (i) FL4BCI-CNN and (ii) FL4BCI-LSTM. Specifically, the architecture of FL4BCI-CNN is designed to match that of FL4BCI-SNN, while the FL4BCI-LSTM network consists of two LSTM layers followed by a single linear layer.
@@
[…]
As aforementioned, both identification accuracy and energy consumption play a crucial role in determining the QoS provided by MSPs and the QoE experienced by users. As a result, we introduce a new metric Weighted System Performance (WSP) that combines identification accuracy and energy consumption. By integrating these factors into this metric and weighting them equally, we can ensure a balanced evaluation of user satisfaction across various system configurations. This metric allows for a comprehensive analysis of the trade-offs between identification accuracy and energy efficiency in service delivery, providing a holistic view of overall system performance.
As illustrated in Fig. 5, FL4BCI-SNN demonstrates WSP improvements of 2.12 and 2.11 times compared to the baseline methods, significantly outperforming both FL4BCI-CNN and FL4BCI-LSTM. This result confirms the superior efficacy of our proposed FL4BCI-SNN model in enhancing WSP. Note that although the constructed WSP metric does not currently integrate with other indicators that influence QoE and QoS, such as resource allocation schemes and wireless communication quality, it is designed to be adaptable and can easily be combined with these factors.
Figure 5: The achievable WSP of various methods.
@@ V Challenges and Future Research Directions
In this section, we discuss the practical deployment challenges faced by Bio-HDT from two perspectives: brain signal processing and SNN-driven FL. Additionally, we explore future research directions based on the aforementioned analysis.
@@ V-A Challenges on Processing Brain Signals
The primary challenge in deploying the Bio-HDT system lies in the accurate and real-time processing of brain signals. Given the system’s reliance on continuous monitoring of cognitive, emotional, and behavioral states, the quality of signal extraction and interpretation is paramount. Brain signals, often noisy and prone to external interference, exhibit significant variability between users. Developing robust signal processing techniques that effectively filter out noise while retaining critical features is essential. Additionally, personalized algorithms may be necessary to account for individual differences, as a generic, one-size-fits-all approach may not be effective across diverse user bases.
Another critical factor is ensuring real-time synchronization between brain activity and HDTs within the future human-centric Metaverse. This synchronization is crucial for maintaining an immersive and responsive user experience. With anticipated advancements in B5G/6G systems, the Bio-HDT framework is prone to overcome existing latency challenges by enabling near-instantaneous data transmission. The advent of 6G technology has the potential to revolutionize Bio-HDT interactions by leveraging ultra-low latency, edge computing, and enhanced data rates to deliver seamless and immersive experiences within the Metaverse. These advancements will enable brain signals to be processed and responded to in near real-time, significantly enhancing the depth and responsiveness of user interactions.
@@ V-B Challenges in Implementing SNN-driven FL
One of the primary challenges in deploying SNN-driven FL within the Bio-HDT system lies in the training complexity of SNNs. Unlike traditional ANNs, which process continuous signals, SNNs rely on discrete spikes of electrical activity to transmit information. This discrete, event-driven nature introduces unique difficulties during training. Conventional optimization techniques, such as gradient-based methods (i.e., backpropagation), are less effective for SNNs due to the discontinuous nature of the spike function. Although several approaches, such as surrogate gradient techniques [12] and ANN-to-SNN conversion [11], have been proposed to address these challenges, further exploration is essential to fully unlock the potential of SNNs in real-world applications.
Another significant challenge is the non-independent and identically distributed nature of brain signal data, primarily caused by neurodiversity among users. Devices such as BCI-driven HMDs process brain signals that vary widely due to each individual’s unique neural architecture and cognitive patterns. This variability leads to significant differences in the data across devices, complicating the aggregation and training of a coherent global model in FL. One potential solution is to design cross-learning algorithms that enable efficient transfer learning, allowing the model to adapt to diverse brain signals and improve generalization across users with varying neural characteristics.

### v3

@@ IV-C Case Study
We demonstrate the efficacy of our framework through comparison simulations in the following. As aforementioned, the Bio-HDT framework is primarily designed to enhance three key factors: personalization, energy efficiency, and privacy preservation. In the human-centric Metaverse, personalization relies heavily on the accurate interpretation of human’s behavior intention and states, ensuring that user interactions are not only immersive but also highly responsive. For MSPs, the ability to accurately identify and interpret these signals is crucial. This capability is fundamental to effectively analyzing user requirements and creating personalized services that meet individual needs. The precision in processing these signals directly influences the quality, relevance, and timeliness of the content delivered. Consequently, it impacts both the QoS provided by the MSPs and the QoE perceived by the users. Additionally, energy efficiency is crucial for enabling sustainable and scalable operations in real-time environments, ensuring that the system can operate continuously without excessive resource consumption. Therefore, we evaluate our scheme’s performance based on the identification accuracy of brain signals and the energy consumption of neural network operations. Note that privacy is not evaluated in this case study because our SNN-driven FL framework effectively addresses privacy concerns by processing data locally. The settings of simulation are as follows:
IV-C1 BCI data statement
@@ IV-C3 FL training and baselines
We utilize the FedAvg algorithm [McMahan2016CommunicationEfficient] to conduct the FL framework. For the learning hyper-parameters, the learning rate, batch size and local epoch are set as $\left\{0.01,64,5\right\}$. Moreover, we conduct performance comparisons between our proposed method, denoted as FL4BCI-SNN, and two baseline methods: (i) FL4BCI-convolutional neural network (CNN) [LeCun2015Deep] and (ii) FL4BCI-long short-term memory (LSTM) [LeCun2015Deep]. The FL4BCI-CNN baseline is constructed using a traditional ANN-based CNN architecture and adopts an identical structural configuration to FL4BCI-SNN, comprising two convolutional layers, each followed by batch normalization, ReLU activation, and pooling, and one final linear layer for classification. The FL4BCI-LSTM baseline consists of two stacked LSTM layers with a single linear output layer, enabling the model to capture temporal dependencies within BCI signals. This consistent architecture design ensures fair comparison between the SNN-based and ANN-based frameworks in terms of depth and parameter scale.
@@ V Challenges and Future Research Directions
In this section, we discuss the Bio-HDT deployment challenges and outlines future research directions.
@@ V-A Challenges on Processing Brain Signals
The primary challenge in deploying the Bio-HDT system lies in the accurate and real-time processing of brain signals. Brain signals, often noisy and prone to external interference, exhibit significant variability among users (i.e., neurodiversity). Developing robust signal processing techniques that effectively filter out noise while retaining critical features is essential. Additionally, personalized algorithms may be necessary to account for individual differences, as a generic, one-size-fits-all approach may not be effective across diverse user bases.
Another critical factor is maintaining real-time synchronization between neural activity and its HDT representation, which is essential for immersion and responsiveness in a human-centric Metaverse. Emerging B5G/6G techniques can reduce end-to-end delay across sensing, inference, and actuation, enabling near-real-time HDT updates. By executing SNN inference and FL aggregation at the edge, the Bio-HDT pipeline can process and respond to EEG streams promptly, thereby deepening interaction fidelity while preserving user experience.
@@ V-B Challenges in Implementing SNN-driven FL
One of the primary challenges in deploying SNN-driven FL within the Bio-HDT system lies in the training complexity of SNNs. Unlike traditional ANNs, which process continuous signals, SNNs rely on discrete spikes of electrical activity to transmit information. This discrete, event-driven nature introduces unique difficulties during training. Conventional optimization techniques, such as gradient-based methods (i.e., backpropagation), are less effective for SNNs due to the discontinuous nature of the spike function. Although several approaches, such as surrogate gradient techniques and ANN-to-SNN conversion [9543525], have been proposed to address these challenges, further exploration is essential to fully unlock the potential of SNNs in real-world applications.
Another significant challenge is the neurodiversity of of brain signals. This variability leads to significant differences in the data across devices, complicating the aggregation and training of a coherent global model in FL. One potential solution is to design cross-learning algorithms that enable efficient transfer learning, allowing the model to adapt to diverse brain signals and improve generalization across users with varying neural characteristics.